OpenStudy (anonymous):

Represent 375 and 4856 in the Babylonian number system

OpenStudy (primeralph):

Now what?

OpenStudy (primeralph):

@robtobey

terenzreignz (terenzreignz):

Babylonian number system? -.-

OpenStudy (anonymous):

i am have trouble understanding how to put these numbers together using their base 60

terenzreignz (terenzreignz):

Too bad these aren't just numbers from 1 to 59, or this'd be ridiculously trivial :D

OpenStudy (primeralph):

Wait, so we have to draw those sticks or whatever?

terenzreignz (terenzreignz):

Unless LaTeX can provide for them :P

OpenStudy (anonymous):

ha, yeah.

ganeshie8 (ganeshie8):

guess its positional number system with out 0, with 60 digits/symbols

OpenStudy (primeralph):

@zonazoo What are you studying by the way?

OpenStudy (primeralph):

So mod 60.

ganeshie8 (ganeshie8):

need to account for missing 0 also not sure

terenzreignz (terenzreignz):

Okay, so... it seems this is just... place value system with base 60

OpenStudy (primeralph):

http://gwydir.demon.co.uk/jo/numbers/babylon/

terenzreignz (terenzreignz):

Let's start with 375.. how many 60's are in 375?

OpenStudy (anonymous):

ok, 6

terenzreignz (terenzreignz):

so... unlike our 'normal' system, where we have a place values as ones, tens, hundreds, etc.. They use a sexagesimal system, they have a ones, 60's, 3600's etc....

terenzreignz (terenzreignz):

So, 375 doesn't reach 3600 yet, so we start counting the number of 60's... That's 6. And what remains?

OpenStudy (anonymous):

15

OpenStudy (primeralph):

@terenzreignz I think the problem is the drawing.

terenzreignz (terenzreignz):

Is it? @zonazoo ?

OpenStudy (anonymous):

the difference from 375, or u mean from 3600

OpenStudy (goformit100):

@terenzreignz did it correctly.

terenzreignz (terenzreignz):

No, I mean, IS the problem just drawing?

OpenStudy (anonymous):

ohh, yes, i dont understand which symbols to use.

OpenStudy (primeralph):

@goformit100 I know @terenzreignz is correct. What I'm saying is, once you get above 9, it's hard to represent other digits. Especially in a case of base 60.

OpenStudy (primeralph):

So you'll need to use the diagrams explicitly.

OpenStudy (goformit100):

ya that is correct.

terenzreignz (terenzreignz):

Here, a place value table... \[\large \left.\begin{matrix}\times & \color{red}{60^2}&\color{red}{60^1}&\color{red}{60^0}\\375 &0&6&15 \\4856 &1 & 20 &56 \end{matrix}\right.\]

OpenStudy (anonymous):

there is an example that is given, if you look at the link with the chart above, there is a drawing of 4 symbols... 1, 57, 46, 40... and these when placed next to each other some how mean, 424,000

terenzreignz (terenzreignz):

Yes... it basically means.... \[\Large 1\times 60^3 + 57\times60^2 + 46\times60 + 40\times 60\]

OpenStudy (anonymous):

okay, i understand your table.

OpenStudy (anonymous):

so your saying for 375, i am going to use the 6th and 15th symbol.

OpenStudy (primeralph):

flgbrgerelrjernermernxpodrnxe That's Babylonian for good night fellas.

terenzreignz (terenzreignz):

Whoops, I meant \[\Large \color{red}{1\times 60^3} + \color{green}{57\times60^2} +\color{blue}{ 46\times60^1} + 40\times 60^0\]

terenzreignz (terenzreignz):

And yes, @zonazoo

OpenStudy (anonymous):

okay... lets just make up some random number of 53,849.... I would have to figure out how many times 3600 goes into this just to get the first symbol

terenzreignz (terenzreignz):

Yes... Unless somehow, 216000 also goes in...

terenzreignz (terenzreignz):

But it doesn't... So... how many times does 3600 go in?

OpenStudy (anonymous):

ummm, looks to be like 13 or 14... maybe 15.... im trying to get this in my head.

terenzreignz (terenzreignz):

14. Don't hurt yourself :D

OpenStudy (anonymous):

that seems like what would happen when u deal with large numbers, i couldnt imagine putting this much effort into figuring out how to write a number... so okay 14 x 60^2 + 57 x 60^1 + 29

terenzreignz (terenzreignz):

Yup :) so in babylonian numerals, it'd be 14th symbol, 57th symbol, 29th symbol :D

terenzreignz (terenzreignz):

We actually use the sexagesimal all the time... literally, when telling time :D 60 seconds, 60 minutes... well, that ends there :P

OpenStudy (anonymous):

ha. well thanks for all your help.

terenzreignz (terenzreignz):

No problem... :)